OpenStudy (anonymous):

numbers that are divisible by 22 and 7

Parth (parthkohli):

\[22 = 11\cdot 2 , \ \ 7 = 7\]So the list starts with \(11\cdot 2 \cdot 7\), then we can have these numbers:\[11\cdot 2^2 \cdot 7 \\ 11 \cdot 2 \cdot 7^2 \\ 11^2 \cdot 2 \cdot 7^2\]Et cetera

OpenStudy (shubhamsrg):

are divisible by 22*7 .Ain't that nice ?

Parth (parthkohli):

Yup, if the numbers are coprime, that applies. ^ :-)

OpenStudy (anonymous):

Actually my real question was if an equilateral triangle square regular hexagon and circle of same perimeter are given which has largest area?

OpenStudy (shubhamsrg):

oh, I see. :)

OpenStudy (shubhamsrg):

Mistakes happen buddy, that is nothing to worry .

OpenStudy (shubhamsrg):

We also misunderstood your query perhaps.

OpenStudy (shubhamsrg):

We all are humans.

OpenStudy (shubhamsrg):

One who doesn't make mistakes is not human.

OpenStudy (shubhamsrg):

amen.

OpenStudy (anonymous):

thank u 4 the support

OpenStudy (shubhamsrg):

My support is always with you buddy.

OpenStudy (shubhamsrg):

You're god's one beautiful creation, how can i dare not to support you!

OpenStudy (shubhamsrg):

You're one of the smarter and wise and what not, persons I've ever seen on OS till date.

OpenStudy (shubhamsrg):

Let us all, together, say thanks, to the almightly, to have created you, to have created us, the human race.

OpenStudy (shubhamsrg):

amen.

Parth (parthkohli):

...lol